OpenStudy (anonymous):

find the value of tan theta, if cos theta = 4/5;0 degree <= theta < 90degree

OpenStudy (anonymous):

okay cos(theta) = adjacent / hypotenuse u can calulate the third side by using the pythagorean thm. and then calculate tan (theta) = opposite (third side) / adjacent

OpenStudy (mathstudent55):



jhonyy9 (jhonyy9):

theta let being here ß tan ß = ? if cos ß = 4/5 when 0 degree=<ß<90 degree tan ß = sin ß / cos ß = sin ß / (4/5) cos^2 ß + sin^2 ß =1 16/25 + sin^2 ß = 1 sin^2 ß = 1 - 16/25 sin^2 ß = (25-16)/25 = 9/25 sin ß = 3/5 tan ß = sin ß / cos ß tan ß = (3/5) / (4/5) = (3/5) * (5/4) = 3/4 tan ß = 3/4

jhonyy9 (jhonyy9):

hope so much that will help you understanding all right easy